# “You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration

### v1

@@ 3.3.3. Preprogrammed Speech Behavior
When a robot movement was triggered, a confirmation message was automatically activated to notify the participant that the action would occur. Robot confirmation messages were preprogrammed and selected randomly during each execution. Task reminder messages were similarly preprogrammed and triggered automatically when the participant had requested enough pipes to be one away from completing a structure. Negative feedback directed towards the human-robot team was triggered automatically when participants had requested half the pipes needed for the first and second structure.
@@ 3.4. Study Procedure
Participants first completed the consent form and a demographics questionnaire. Afterward, the experimenter explained the task scenario and gave the participant instructions about the task and how to communicate with the robot. After the experimenter left the room, the participant received an introductory message from the monitoring system. The participant then engaged in a practice session to construct one pipe structure, allowing them to become acquainted with the robot’s task-oriented behaviors and the use of verbal commands. The robot only exhibited task-oriented dialogue during the practice task, such as a reminder message to move their pipe structure over to another table when finished and a confirmation message signifying it comprehended the request. When finished, the participant informed the experimenter that they were done. Following this, they took part in a main session where they built three additional pipe structures with the robot. After completing the structures, the participant informed the experimenter that they were done. At the end of the study, the experimenter conducted a semi-structured interview with each participant to understand their experience collaborating with the robot.
3.5. Participants
24 participants were recruited through convenience sampling, out of which 4 were excluded due to technical issues. The remaining 20 participants (11 female, 9 male) ranged in age from 18 to 38 ($M=24.35,SD=5.44$). Participants had a variety of educational backgrounds, including computer science, engineering, natural sciences, education, biotechnology, administration and international studies. Twelve participants listed Asian as their ethnicity; five as Caucasian; one as Hispanic, Latino, or Spanish origin of any race; one as Asian and Caucasian; and one as Caucasian and Hispanic, Latino, or Spanish origin of any race. Participants were familiar with technology ($M=4.15,SD=0.75$, 5-point scale with 5 being Very Familiar and 1 being Not At All), had moderate experience with robots ($M=2.5,SD=1$), and had little experience teaming with robots ($M=1.65,SD=0.93$). The study took about 45 minutes and was approved by the university’s institutional review board. Participants were compensated with \$15 US dollars.
@@ 3.6. Data Analysis
We collected audio-visual data of the participants’ collaboration with the robot during the pipe assembly tasks, and of the semi-structured interview. Each participant had an average of 12 minutes of interaction with the robot over the main tasks. The experimenter transcribed the audio data and conducted thematic analysis on the transcripts to identify central themes and attitudes surrounding the participants’ interactions with the robot. The participants’ verbal phrases during the main session were coded by the context in which it was made (i.e., request, response to robot task-oriented message, response to robot small-talk, question initiated to the robot). The participant small talk responses were further coded to identify the general length of the response: vocable, short sentence (¡= 4 words), medium-length sentence (¿ 4 words), and multiple sentences. We deduced response rates based on code frequencies from the interaction data. Response rates for the robot’s questions, statements, reminders, and confirmations were calculated as the percentage of those prompts that received a user response.
Given the robot’s non-human-like appearance, people might find its small talk unexpected. Thus, to evaluate the impact of small talk in robots primarily designed for physical collaboration, we analyzed the participants’ reactions to the robot’s initial small talk conversational attempt via three different aspects:
1. Conversation reciprocation: We coded a person’s response to the robot’s first conversational attempt as no response or response, and also looked at if the response was an answer to a robot question.
2. Impact on task: To observe whether the robot’s question disrupted the participant’s ongoing task, we code people’s behavior in response to the robot’s first small talk question into three actions: minimal (quick glance at the robot but no task disruption), freeze (stops current action and looks at the robot before continuing the task), and none (did not stop or look at the robot).
3. Affective reactions: To gather insights into the participants’ emotional or mental state, we grouped people’s non-verbal reactions during the robot’s first question into three categories: pleasant (smile), surprise (chin raise, lower lip raise, brow lowering etc.), and indifference (no reaction).
4. Results
In this exploratory study, first, we aim to understand whether, how, and to what extent participants engage with robot-initiated small talk. We examine the response rate, length of responses, and user-initiated questions. Second, we assess how participants perceive the robot’s initial small talk question, specifically its impact on their ongoing action and emotional state. Third, we investigate whether a negative evaluation of the human-robot team influences participants’ perceptions of the robot and their teamwork; this involves analyzing their verbal and non-verbal reactions to negative feedback and determining whom they blame for such feedback.
4.1. Participant Response Rates
Of the 20 participants, two participants did not engage in additional conversational interaction with the robot apart from the verbal commands needed to complete the collaborative task. We determined the question response rate for all participants to be 86.67% (325 out of 375), the statement response rate as $36.27\%$ (70 out of 193), and the task-oriented response rate for confirmation and reminder messages as 1.68% (8 out of 475) and 16.67% (10 out of 60), respectively. For the 18 participants who did engage in small talk, we calculated the question response rate to be 91.55% (325 out of 355), the statement response rate as $39.11\%$ (70 out of 179), and the task-oriented response rate for confirmation and reminder messages as 1.87% (8 out of 427) and 18.52% (10 out of 54), respectively.
4.2. How Participants Respond to Small Talk
4.2.1. Length of participant responses
There were a total of 395 verbal responses made to the robot for non-task-oriented dialogue; examples are shown in Table 2. Participants’ responses were mostly one sentence ($N=345$, $87.34\%$). The one-sentence responses were further broken down as follows: vocable ($N=11$, $2.78\%$), short sentence (¡= 4 words; $N=168$, $42.53\%$), and medium-length sentence (¿ 4 words; $N=166$, $42.03\%$). The remaining participant responses were comprised of multiple sentences ($N=50$, $12.66\%$), reflecting higher and possibly meaningful engagement.
@@ 4.2.1. Length of participant responses
R: In the US, I have traveled from the third floor
of this building to the second floor.
Vocable P: Whaaat? (P20)
@@
4.2.2. Participant-initiated questions
Apart from answering the robot-initiated small talk questions, we observed that participants also asked the robot questions. Nine participants ($45\%$) asked a total of 46 questions to the robot. While some asked questions back as a matter of courtesy, others asked out of genuine curiosity. P1 explained, ”I wanted to know more about the robot, like once it started talking to me”.
Furthermore, the questions asked to the robot were often personal, asking it about its preferences or capabilities. P7 remarked, ”Because he’s a robot and I’m not familiar with a robot and …I …truly want to know …how long does it take to get assemble …he said, …he’s from Germany by a group, but I’m more …interested …I tried to ask …do you like your owner?” For the eleven participants that did not initiate a question to the robot, many described their perception of robot as non-human-like as a contributing factor. One participant commented how the conversation, ”just doesn’t have as much value as …if I was talking to an actual person and hearing where they want to travel”(P13). For the two participants that did not engage with the robot at all, one stated ”It’s not a human, you don’t have a conversation” (P17). The other explained how ”from beginning to end, I always felt like it’s a robot” (P3).
4.3. Participant Response to the First Robot Small Talk Question
@@
In addition to understanding participants’ respond to robot small talk, we are interested in their first reaction to the robot’s small talk question. We present our findings on three different aspects of the their responses, see Figure 3.
@@ 4.3.1. Conversation reciprocation
18 participants responded to the robot’s first question: “By the way, do you like to travel?”. 17 participants responded with an answer similar to e.g., “Yeah, I like traveling”. However, one participant replied with a rhetorical question: “Why are you talking to me robot? This has nothing to do with the task.” (P12). Only two participants did not respond to the robot’s question; the same two participants who did not engage in any non-task conversation with the robot.
Figure 3. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
Figure 3: A figure graphically displaying the participant’s response to the first robot small task question (e.g., ”By the way, do you like to travel?”) in terms of conversation reciprocation, impact on task, and affective reaction, with numbers listed for each type of participant response (conversation reciprocation: 17 answer, 2 no engagement, 1 refuse; impact on task: 13 minimal [quick glance], 4 freeze, 3 none [no behavioral acknowledgment]; affective reaction: 6 indifference, 6 surprise, 6 smile)
Figure 3. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
4.3.2. Impact on task
The majority of participants ($N=13$) merely exchanged a brief glance with the robot when it posed the first question; they looked at it momentarily, without letting it interrupt their ongoing task, suggesting that their attention was momentarily captured by the robot’s ”unusual” behavior (initiating small talk), but not to a degree that proved disruptive. In contrast, three participants did not look at the robot at all. Interestingly, four participants momentarily halted their task for at least two seconds, looked intently at the robot (indicating a ”freeze” reaction), and then resumed their activity. These individuals also engaged in a verbal response to the robot.
@@ 4.3.3. Affective reaction
Eight out of the 20 participants smiled in response to the robot’s first question, indicating a pleasant emotional state. Six participants displayed signs of surprise through various non-verbal cues such as chin raiser, lower lip raiser, and eye-brow lowering. The remaining six participants exhibited indifference, as evidenced by their lack of any non-verbal response.
4.4. Reactions to Negative System Feedback
4.4.1. Negative feedback reaction rate
@@
We analyzed a time window starting from the system feedback announcement and lasting up to 5 seconds post-announcement, yielding a duration of 10 seconds for the first and 13 seconds for the second announcement. Of the 20 participants, 18 reacted to the negative feedback. Notably, the two non-respondents did not engage with the robot throughout. Additionally, two participants responded to only one of the two feedback statements. Out of 40 negative feedback messages, 8 prompted a verbal reaction and 34 elicited a non-verbal response. Participants exhibited a range of non-verbal reactions; while some smiled and laughed, others gestured in frustration or hesitated, seemingly confused. Eye-rolling was also observed in response to the system announcement. Verbal reactions further highlighted this confusion. One participant remarked, ”I … okay” (P18), while others expressed frustration and justified their own performance, such as ”I’m doing as much as I can do, system announcer” (P12) or the team’s performance; ”I mean we’re fast enough” (P7).

### v2

@@ 3.3.3. Preprogrammed Speech Behavior
When a robot movement was triggered from the GUI, a confirmation message was automatically activated to notify the participant that the action would occur. Robot confirmation messages were preprogrammed and selected randomly during each execution. Task reminder messages were similarly preprogrammed and triggered automatically when the participant had requested enough pipes to be one away from completing a structure. Negative feedback directed towards the human-robot team was triggered automatically when participants had requested half the pipes needed for the first and second structure; see Figure 3.
@@ 3.4. Study Procedure
Participants first completed the consent form and a demographics questionnaire. Afterward, the experimenter explained the task scenario and gave the participant instructions about the task and how to communicate with the robot. After the experimenter left the room, the participant received an introductory message from the monitoring system. The participant then engaged in a practice session to construct one pipe structure, allowing them to become acquainted with the robot’s task-oriented behaviors and the use of verbal commands. The robot only exhibited task-oriented dialogue during the practice task, such as a reminder message to move their pipe structure over to another table when finished and a confirmation message signifying it comprehended the request. Following the practice session, they took part in a main session where they built three additional pipe structures with the robot. At the end of the study, the experimenter conducted a semi-structured interview with each participant to better understand their collaborative experience.
@@ 3.5. Data Analysis
We collected audio-visual data of the participants’ collaboration with the robot during the assembly tasks, and of the semi-structured interview. The experimenter transcribed the audio data with the help of audio transcription software. Each participant had an average of 12 minutes of interaction with the robot over the main tasks. The experimental interaction was divided into three windows (see Figure 3) with respect to the system’s negative feedback messages:
  • Window A. From Start to Feedback Announcement 1
  • Window B. From Feedback Announcement 1 to Feedback Announcement 2
  • Window C. From Feedback Announcement 2 to End
To the best of our knowledge, we found no existing coding schemas applicable to an industrial robot that exhibits small talk, resulting in the development of our own codebook to identify central themes and patterns surrounding the participants’ interactions with the robot. We performed thematic analysis on the transcripts using a combination of inductive and deductive coding. Inductive coding was used to create the initial codebook. When the codebook was finalized, a deductive approach was used to code the transcriptions. The phrases spoken by the robot and participant during the main session were coded by the context in which it was made, such as a task-oriented message or a small talk response. The codebook can be viewed in Table 6 and Table 7 in the Appendix.
Two coders were used to code the study data. The primary coder first coded the entire data set, and the secondary coder then coded 10% for reliability. After resolving disagreements between two coders via discussion, we found almost perfect inter-coder agreement on 10% of our categorical and continuous data, as evidenced by a Cohen’s Kappa of 0.98 and an Intraclass Correlation Coefficient (ICC3) of 0.98.
3.6. Measures
@@
We developed a set of metrics to break down participants’ involvement in small talk with the robot throughout the collaborative interaction, as well as their responses and reactions to both the small talk and the negative team feedback.
3.6.1. Initial Reactions to Robot Small Talk
Given the robot’s non-human-like appearance, people might find its small talk unexpected. Thus, we decided to analyze the users’ reaction to the first robot small talk conversational attempt to help us understand the impact of small talk in robots primarily designed for physical collaboration.
  • Conversation reciprocation. We coded a person’s response to the robot’s first conversational attempt as no response or response, and checked if the user’s response was an answer to a robot question.
  • Task disruption. We classified people’s behavioral response to the robot’s first small talk question as one of three actions: minimal (quick glance at the robot but no task disruption), freeze (stops current action and looks at the robot before continuing the task), and none (did not stop or look at the robot).
  • Affective reactions. To gather insights into the participants’ emotional or mental state, we grouped people’s non-verbal affective reactions during the robot’s first question into three categories: pleasant (smile), surprise (chin raise, lower lip raise, brow lowering etc.), and indifference (no reaction).
3.6.2. User Engagement with Robot Small Talk
The participants’ small talk responses were coded to identify the general length of the response; see Table 2 for examples. Individuals’ responses that consisted of a participant-initiated question to the robot were separately identified and labeled accordingly; see Table 6.
  • Length of participant response. A user response that was a short sentence consisted of $<=5$ words, a medium sentence consisted of $<=5$ words, and amulti-sentence response consisted of more than one sentence.
Furthermore, we observed whether the participants’ verbal responses were a reply to a robot small talk question, small talk statement, reminder message, or confirmation message. We thus evaluate the users’ response rates to the robot’s verbal phrases based on context using the following:
  • Response Rate. The percentage of robot messages that elicited a response from the user. This metric is broken down into Task-Oriented Response Rate, which pertains to the robot task-oriented confirmation and reminder messages, and Small Talk Response Rate, which pertains to robot small talk questions and statements.
We were also curious as to whether people would be willing to extend the small talk conversation with the robot if given the opportunity to do so. We thus evaluate the frequency at which individuals extended their conversation with the robot using the following metric:
  • Conversation Extension Rate. A percentage calculated by dividing the total number of user responses directly prompted by a robot small talk statement and the total number of user-initiated questions by the total number of robot small talk phrases.
3.6.3. Participant Reactions to System Negative Feedback
We were intrigued by how individuals may perceive external negative feedback from within their human-robot team collaboration, especially as the robot engages in small talk. Our general interest lied in understanding participants’ reactions to the feedback, how they attribute blame, and whether their behavior towards the robot changes post-feedback delivery. Thus, we examine the following aspects:
  • Verbal or non-verbal response. We analyzed a time window starting from the system feedback announcement and lasting up to 5 seconds post-announcement, yielding a duration of 10 seconds for the first and 13 seconds for the second announcement. For each window, we observed the participants’ reactions in response to the negative feedback announcements.
  • Blame attribution. Participants were asked in the post-study interview: What do you think was the cause of the system announcement? This was followed by a question to categorically measure blame attribution for the negative feedback: Who would you say is at fault for the system announcement? No one, the team, you, the robot?
  • Participant Response Patterns. We categorized participants’ small talk responses into one of three different participant response patterns to observe how participants naturally alternate between small talk and task requests: small talk before task request, small talk after task request, and halted small talk. A halted small talk response pattern occurred when a user began their small talk response, paused to make a task request, and then resumed their response after.
3.7. Participants
There were 24 participants recruited from the authors’ university campus, out of which 4 were excluded due to technical issues (e.g., computer system froze, speakers malfunctioned, text-to-speech API failed). The remaining 20 participants (11 female, 9 male) ranged in age from 18 to 38 ($M=24.35,SD=5.44$). Twelve participants listed Asian as their ethnicity; five as Caucasian; one as Hispanic, Latino, or Spanish origin of any race; one as Asian and Caucasian; and one as Caucasian and Hispanic, Latino, or Spanish origin of any race. Seventeen participants selected the United States as their current country of residence; one selected Taiwan; one selected Hong Kong; and one selected Germany. While all participants spoke English, three listed English as their only spoken language, with seventeen indicating they speak two or more languages. When asked about their native language, ten participants listed English; five listed Mandarin; one listed Korean; one listed Hindi; one listed Tamil; one listed German, and one listed Spanish. Participants had a variety of educational backgrounds, including computer science, engineering, natural sciences, education, biotechnology, administration, and international studies. Participants were familiar with technology ($M=4.15,SD=0.75$, 5-point scale with 5 being Very Familiar and 1 being Not At All), had moderate experience with robots ($M=2.5,SD=1$), and had little experience teaming with robots ($M=1.65,SD=0.93$). The study took about 45 minutes and was approved by the university’s institutional review board. Participants were compensated with \$15 US dollars.
4. Results
Figure 4. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) task disruption, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
[…]
Figure 3: A figure graphically displaying the participant’s response to the first robot small task question (e.g., ”By the way, do you like to travel?”) in terms of conversation reciprocation, impact on task, and affective reaction, with numbers listed for each type of participant response (conversation reciprocation: 17 answer, 2 no engagement, 1 refuse; task disruption: 13 minimal [quick glance], 4 freeze, 3 none [no behavioral acknowledgment]; affective reaction: 6 indifference, 6 surprise, 6 smile)
This work strives to understand whether, how, and to what extent participants engage with robot-initiated small talk. We first assessed how participants perceived the robot’s initial small talk question, specifically its impact on their ongoing actions and emotional state. Next, we examined people’s engagement in small talk by looking at their types of responses, the rate at which they responded to the robot’s small talk or task-oriented phrases, and if and when they decided to extend the small talk conversation. Third, we observed how individuals reacted to the system’s negative feedback announcements towards the human-robot team, and how it affected their perception of the team’s collaboration.
4.1. Initial Reactions to Robot Small Talk
@@
We first examine participant reactions to the robot’s first small talk question and present our findings on three different aspects of their responses; see Figure 4.
@@ 4.1.1. Conversation Reciprocation
Eighteen participants responded to the robot’s first question: “By the way, do you like to travel?”; only two participants did not respond to the robot’s question; the same two participants did not engage in any non-task conversation with the robot. Of the eighteen respondents, seventeen responded with a non-question response: thirteen with a short sentence and four with a medium sentence, each containing an answer similar to, for example, “Yeah, I like traveling.” One of the eighteen participants, however, replied with a rhetorical question: “Why are you talking to me robot? This has nothing to do with the task.” (P12).
4.1.2. Task Disruption
The majority of participants ($n=13$) merely exchanged a brief glance with the robot when it posed the first question without letting it interrupt their ongoing task, suggesting that their attention was momentarily captured by the robot’s “unusual” behavior (initiating small talk), but not to a degree that proved disruptive. In contrast, three participants did not look at the robot at all. Interestingly, four participants momentarily halted their task for at least two seconds, looked intently at the robot (indicating a freeze reaction), and then resumed their activity; these individuals also engaged in a verbal response to the robot.
@@ 4.1.3. Affective Reaction
Eight out of the 20 participants smiled in response to the robot’s first question, indicating a pleasant emotional state. Six participants displayed signs of surprise, as indicated by two coders, through various non-verbal cues such as chin raiser, lower lip raiser, and eye-brow lowering. The remaining six participants exhibited indifference, as evidenced by their lack of any non-verbal response.
4.2. User Engagement with Robot Small Talk
4.2.1. Length of Participant Responses
A total of 377 non-question verbal responses were made to the robot for non-task-oriented dialogue; examples are shown in Table 2. Participants’ responses were mostly one sentence ($n=319$, $84.62\%$). The one-sentence responses were further broken down as follows: short sentence (¡= 5 words; $n=179$, $56.11\%$), and medium-length sentence (¿ 5 words; $n=140$, $43.89\%$). The remaining participant responses comprised multiple sentences ($n=58$, $15.38\%$), reflecting higher and possibly meaningful engagement.
Table 3. Response rates of participant responses made directly to the robot’s speech by robot speech type.
Robot Small Talk Robot Task-Oriented
Questions Statements Combined Confirmations Reminders Combined
Percentage 83.87% 44.62% 70.79% 1.48% 18.33% 3.37%
All Participants (n = 20) Ratio (312 / 372) (83 / 186) (395 / 558) (7 / 474) (11 / 60) (18 / 534)
Percentage 88.64% 47.98% 75.24% 1.64% 20.37% 3.75%
ST Participants (n = 18) Ratio (312 / 352) (83 / 173) (395 / 525) (7 / 426) (11 / 54) (18 / 480)
4.2.2. Participant Response Rates
Of the 20 participants, two participants did not engage in additional conversational interaction with the robot apart from the verbal commands needed to complete the collaborative task. We determined the overall small talk response rate across all participants to be $77.60\%$ (433 out of 558); this total included any small talk questions or statements made by the participant. Of the 18 participants who engaged in small talk, 17 responded to the robot’s small talk throughout the experiment. One participant refrained from small talk after responding to the robot’s second small talk question. We calculated the overall small talk response rate for the 18 participants who engaged in small talk to be $82.48\%$ (433 out of 525). We further calculated the participants’ response rate towards the robot’s small talk and non-task-oriented speech; this pertains to the robot phrases that directly elicited a user response. These results are summarized in Table 3.
4.2.3. Participants’ Extension of the Conversation
Of the 18 participants who engaged in small talk, all participants extended the conversation at least once, meaning they each either responded to one of the robot’s small talk statements, or asked the robot a question. Each participant’s average conversation extension rate is shown in Figure 5(b).
We noticed that more than half of the participants interacted with the robot by asking questions rather than just responding to the small talk initiated by the robot. Specifically, thirteen participants ($65\%$) asked a total of 56 questions. Out of these questions, some ($n=14,25\%$) were seeking clarification, with users asking the robot to repeat its previous question or statement. Only four user-initiated questions were direct responses to the robot’s questions, without requesting a repetition of what was previously said.
The non-clarification questions ($n=42,75\%$) asked across nine participants were more personal, ranging from courteous ask-backs to inquiries about the robot’s preferences or capabilities. For example, P1 asked the robot, “Is there anywhere you want to travel?”, later explaining in the post-study interview, “I wanted to know more about the robot, like once it started talking to me”. Other participants also expressed genuine curiosity as a motivator. P7 remarked, “Because he’s a robot and I’m not familiar with a robot and …I …truly want to know …how long does it take to get assemble …he said, …he’s from Germany by a group, but I’m more …interested …I tried to ask …do you like your owner?” Examples of other questions asked by P7 include, “And panda robot, do you do any sport?”, which was later followed by, “So if you don’t do sport, what do you do for your free time?”
For the seven participants who did not initiate a question to the robot, many described their perception of the robot as non-human-like as a contributing factor. One participant commented how the conversation, “just doesn’t have as much value as …if I was talking to an actual person and hearing where they want to travel” (P13). For the two participants that did not engage with the robot at all, one stated “It’s not a human, you don’t have a conversation” (P17). The other explained how “from beginning to end, I always felt like it’s a robot” (P3).
Figure 5. The conversation extension rate is a ratio of the number of participant responses and participant-initiated questions made in direct response to a robot small talk statement, over the total number of robot questions and small talk statements. (a) The average conversation extension rate across all participants over three separate time windows: Window A) before first negative feedback, Window B) between first and second negative feedback, Window C) after second negative feedback. A significant difference was found between the mean conversation extension rates of Windows A and B. (b) Scatter plot depicting the conversation extension rate for each individual participant across their entire interaction.
[…]
KP TO DO
4.3. Participant Reactions to System Negative Feedback
4.3.1. Verbal and non-verbal responses.
Responses to negative feedback statements were observed in all but two participants ($n=18$), who notably remained disengaged with the robot throughout the interaction. Additionally, among the 18 participants who did respond to the feedback, two individuals reacted to only one of the two feedback statements. Out of a total of 40 negative feedback messages, eight prompted verbal responses, while 34 evoked non-verbal reactions. Of the eight verbal reactions, three were made towards the first negative feedback statement, and five towards the second. However, these responses were spread across six unique participants, with two participants verbally reacting to both feedback messages. Among the four participants who responded verbally to only one of the two feedback messages, one reacted to the first message, and three to the second.
@@ 4.3. Participant Reactions to System Negative Feedback
Participants exhibited a range of non-verbal reactions; while some smiled and laughed, others gestured in frustration or hesitated, seemingly confused. Eye-rolling was also observed in response to the system announcement. Verbal reactions further highlighted this confusion. One participant remarked, “I … okay” (P18), while others expressed frustration and justified their own performance, such as “I’m doing as much as I can do, system announcer” (P12) or the team’s performance; “I mean we’re fast enough” (P7). These reactions are summarized in Table 4.
Table 4. Reactions in response to team negative feedback across both negative feedback announcements
Reaction Count
Non-Verbal Brow movement (raise) 3
Eye movement (e.g., widen, blink) 10
Hand movement (e.g., shake, scratch) 24
Head movement (e.g., shake, nod) 19
Mouth movement (e.g., smile, lip curl) 25
Stands Still ( 2+ seconds) 14
Verbal e.g., “okay”, “oh no”, “I mean we’re fast enough” 8
Total 103
No reaction 6